OpenStudy (anonymous):

FAN+MEDAL A group of 10 friends had lunch together at a restaurant. The meal cost a total of $99.50, including a 15% tip. How much was the total bill for lunch before tip?

OpenStudy (anonymous):

there are no choices

OpenStudy (anonymous):

\[x+.15x=99.5\] \[x(1+0.15)=99.5\] \[x=\frac{ 99.5 }{ 1.15 }\] x=86.5217

OpenStudy (anonymous):

84.57

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

i redid it and the answer is total bill for lunch before the tip was $86.52 for sure

OpenStudy (anonymous):

so charlie was right

OpenStudy (anonymous):

x = total bill before the tip 0.15x = tip x + 0.15x = 99.50 1.15x = 99.50 x = 99.50 / 1.15 x = 86.5217391 x = $86.52